OpenStudy (anonymous):

What is the length of the conjugate axis? ((y-2)^2/16)-((x+1)^2/144)=1

OpenStudy (anonymous):

im not sure if this is a vertical hyperbola if sooo then i know i need to use 2b to get the length

OpenStudy (anonymous):

but what do i do with 2b ....?

OpenStudy (sidsiddhartha):

it of the form \[\frac{ (y-\beta)^2 }{ a^2}-\frac{ (x-\alpha)^2 }{ b^2 }=1\] so length of conjugate axis will be\[\large 2b\]

OpenStudy (anonymous):

would b = 4 thennn ???....

OpenStudy (sidsiddhartha):

no compare those equations u'll get \[\large b^2=144\rightarrow b=12\]

OpenStudy (anonymous):

2b = 2(24) = 48 ???

OpenStudy (sidsiddhartha):

how if b=12 then 2b=2*12=24

OpenStudy (anonymous):

OPPPS my bad idk what was that lol

OpenStudy (anonymous):

sorry lol mmmmhhhh okkk i understand now

OpenStudy (anonymous):

thank you n_n @sidsiddhartha